6/13 of the athletes could complete the obstacle course. If 14 of the athletes could not complete the course, how many could complete it?
Answer by vleith(2517)

if 6/13 could complete it, then 1 - (6/13) could not.
7/13 cannot complete it.
7/13 = 14/x
x = (13*14)/7 = 26 . So
26 tried, 14 failed to complete it. leaving 12 to complete it